### v1

@@ 2 Corner excitations
The set of localized potentials $\mathcal{U}^{(t)}$ is a linear subspace of $\mathbb{R}^{J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}}$. The following result gives its complement. The following lemma is simply a rewriting of the governing equation in (1.1) for localized solution in $\mathcal{U}^{(t)}$.
@@
Fix any $\mathbf{u}\in\mathcal{U}^{(t)}$ and extend it onto $D\cup\partial D$ by zero (still denoted by $\mathbf{u}$). By the definition of potential solutions, for $p\in L_{t+1}^{\mathcal{S}}$, we have that $\sum_{q\in\mathcal{N}(p)}\gamma_{pq}(\mathbf{u}_{q}-\mathbf{u}_{p})=0$, or equivalently, that $\mathbf{v}_{p}\cdot\mathbf{u}=0$. Since $\operatorname{supp}(\mathbf{u})=L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}$, we may have $\mathbf{v}_{p}\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}\cdot\mathbf{u}=0$. Subsequently, we have
@@
Conversely, suppose $\mathbf{u}\in\left(\operatorname{span}_{p\in L_{t+1}^{\mathcal{S}}}\left\{\mathbf{v}_{p}\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}\right\}\right)^{\perp}$. Then, for all $p\in L_{t+1}^{\mathcal{S}}$,
$\mathbf{v}_{p}\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}\cdot\mathbf{u}=0,$
After extending $\mathbf{u}$ onto $D\cup\partial D$ by zero, the above equations imply that, conversely, we have $\Delta_{\gamma}\mathbf{u}_{p}=0$ for all $p\in D$, where $\Delta_{\gamma}$ is the graph Laplacian. Therefore,
$\mathbf{u}=S_{\gamma}(\mathbf{u}|_{\partial D})\in\mathcal{U}^{(t)},$
and so
$\mathcal{U}^{(t)}\subseteq\left(\operatorname{span}\left\{\mathbf{v}_{p}\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}:p\in L_{t+1}^{\mathcal{S}}\right\}\right)^{\perp}.$
Combining both inclusions yields the desired orthogonal decomposition. ∎
@@
The definition of excitations and their properties do not rely on the explicit grid graph structure. Whenever one can divide the graph with boundary nodes into two connected parts, one can always define such localized space, and the respective properties remain valid.
@@ 3 Properties of lattice graphs
Note that the unique determination of the conductivity $\gamma$ from the DtN matrix $\Lambda$ is not necessarily valid for a general graph structure. In fact, under the given definitions of $D$ and $\partial D$, we choose the edge set to be complete in $D$ – that is, for every different $p$ and $q$ in $D$, there exists an edge $pq$ in $E$ – then the number of undetermined conductivity parameters, scaling as $O(n^{2d})$, surpasses the size of DtN matrix, which grows as $(2dn^{d-1})^{2}$, for large $n$. This naturally precludes the possibility of unique determination for graph structures, which is characterized by excessive connectivity (i.e. too many edge connections). Thus, the proof of the uniqueness must contain results based on the lattice graph structure.
In this section, we identify several fundamental properties of lattice graphs that may ensure the unique determination of the conductivity $\gamma$. For each integer $t$, consider the subgraph
@@
where the edge set $E^{\prime(t)}$ consists of all edges connecting nodes in $L_{t-1}^{\mathcal{S}}$ to nodes in $\bar{D}$. Associated to $G^{\prime(t)}$ is the DtN map $\Lambda_{\gamma}^{\prime}$, acting on the boundary potentials. Then we define the operator
@@
which maps potentials prescribed on $L_{t}$ to the induced currents on $J_{t-1}^{\mathcal{S}}$.
The next lemma provides the key tool, i.e. to control the potential in the interior domain $L_{t}^{\mathcal{S}}$, using the potential and current taken on $J_{t-1}^{\mathcal{S}}$.
@@
Let $d\leq t\leq dn$ and suppose $\gamma>0$ is a conductivity on the lattice subgraph $G^{\prime(t)}$. The mixed boundary condition problem for $\mathbf{u}\in\mathbb{R}^{L_{t}^{\mathcal{S}}}$,
@@
has only trivial solution, i.e., $\mathbf{u}=\mathbf{0}$.
@@
$H_{l}=(J_{t-1}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}})\cap\{(x_{i})_{i}:x_{1}=l\},\quad H_{l}^{\mathcal{S}}=\bigcup_{i=1}^{l}H_{i}.$
Let $\mathbf{e}_{i}=(0,\ldots,0,{1},0,\ldots,0)\in\mathbb{R}^{d}$ be the $j$-th Cartesian coordinate vector. Note that for each $p\in H_{l}\setminus\partial D$ with $l\geq 1$, the node $q=p-\mathbf{e}_{1}$ is connected to exactly one node that lies outside $H_{l-1}^{\mathcal{S}}$, namely $p$ itself. The base case $\mathbf{u}|_{H_{0}}=\mathbf{0}$ holds by assumption. Now assume that $\mathbf{u}|_{H_{l-1}}=\mathbf{0}$. For each $p\in H_{l}\setminus\partial D$, let $q=p-\mathbf{e}_{1}$. Then there holds
@@
The validity of Lemma (3.1) relies on the graph structure. However, the next lemma is independent of the graph structure.
@@
For every integer $t$ with $d\leq t\leq dn$ and any strictly positive conductivity $\gamma>0$, the following statements are equivalent:
@@
By interpreting the vector $\mathbf{c}$ as a potential supported on $G^{\prime(t)}$ with the zero boundary condition $\mathbf{c}|_{J_{t-1}^{\mathcal{S}}}=\mathbf{0}$, the relation (3.3) is equivalent to the uniqueness in (i):
@@
The injectivity of $T_{2}^{\prime(t)}$ follows directly from Lemmas 3.2 and 3.1. By symmetry, the operator $T_{2}^{(t)}$ is also injective. ∎
@@
where the set $M_{p}:=\mathcal{N}(p)\cap\bigl(J_{t-1}^{\mathcal{S}}\cup L_{t-1}^{\mathcal{S}}\bigr)$ is the neighbours of $p$ within the preceding layer, and the new edge set $E_{p}^{\prime\prime(t)}$ is given by
@@
Note that if $p$ is adjacent to the boundary, isolated boundary nodes may arise. The boundary current on such nodes is set to zero. Fig. 3 gives a schematic illustration of the removal of the node $p$ on the slice $t$.
@@
Figure 3: Node deletion for dimension $d=3$. $L_{t}$ (blue), $L_{t+1}$ (black) and $K_{t}^{-}$ (orange) for $t\leq n+d-1$.
The following lemma establishes the unique recovery of the potential on the reduced vertex set $L_{t}^{\mathcal{S}}\setminus\{p\}$ from the boundary measurements on $J_{t-1}^{\mathcal{S}}$, i.e., the uniqueness of interior potential recovery on the grids $G^{\prime\prime(t)}$.
@@
Let $d\leq t\leq dn$ and fix any node $p\in L_{t}$. Suppose the conductivity $\gamma>0$ is distributed on the subgraph $G_{p}^{\prime\prime(t)}$ defined in (3.4). Consider the boundary value problem:
$\left\{\begin{aligned} \Delta_{\gamma}\mathbf{u}&=\mathbf{0}\quad\text{in }L_{t-1}^{\mathcal{S}}\setminus M_{p},\\ \mathbf{u}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}},\\ D_{\gamma}\mathbf{u}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}},\end{aligned}\right$. (3.5)
@@
Let the deleted node $p$ be located at $(p_{i})_{i}$. We can recover the potential $\mathbf{u}$ in the region $\{(x_{i})\in D:x_{1}<p_{1}\}$ using the same argument in Lemma 3.1. By changing the inductive direction from $\mathbf{e}_{1}$ to $\mathbf{e}_{j}$ with $j>1$, $\mathbf{u}$ can also be recovered in the region
$\{(x_{i})_{i}:\exists i\in\{1,\ldots,d\}\text{ such that }x_{i}<p_{i}\},$
which together with $L_{t}^{\mathcal{S}}\setminus\{p\}$ completes the recovery. ∎
This result yields an important relation among the vectors $\mathbf{v}_{q}$ associated with nodes in the layers $L_{t}^{\mathcal{S}}$, given in the following corollary.
@@
The vectors $\mathbf{v}_{q}$ satisfy the following properties:
@@
    The collection of vectors $\{\mathbf{v}_{q}|_{L_{t-1}^{\mathcal{S}}\setminus M_{p}}\}_{q\in L_{t}^{\mathcal{S}}\setminus\{p\}}$ is linearly independent.
@@
    If $\mathbf{w}\in\operatorname{span}\{\mathbf{v}_{q}|_{L_{t-1}^{\mathcal{S}}}\}_{q\in L_{t}^{\mathcal{S}}}$ is supported on $M_{p}$, then there exists $\alpha\in\mathbb{R}$ such that
    $\mathbf{w}=\alpha\mathbf{v}_{p}|_{L_{t-1}^{\mathcal{S}}}.$
@@
The assertion (i) follows from the argument for the equivalence (i) $\Leftrightarrow$ (ii) in Lemmas 3.2 and $3.$ To see assertion (ii), suppose that $\mathbf{w}=\sum_{q\in L_{t}^{\mathcal{S}}}c_{q}\mathbf{v}_{q}|_{L_{t-1}^{\mathcal{S}}}$ is supported on $M_{p}$. By restricting $\mathbf{w}$ to $L_{t-1}^{\mathcal{S}}\setminus M_{p}$, we have
$\mathbf{0}=\sum_{q\in L_{t}^{\mathcal{S}}\setminus\{p\}}c_{q}\mathbf{v}_{q}|_{L_{t-1}^{\mathcal{S}}\setminus M_{p}}.$
By part (i), it follows that all coefficients $c_{q}=0$ for $q\neq p$. Hence, $\mathbf{w}=\alpha\mathbf{v}_{p}|_{L_{t-1}^{\mathcal{S}}}$ for some $\alpha\in\mathbb{R}$. ∎
Finally, we state an elementary observation on connectivity:
@@
Notably, there are no edges connecting nodes within $L_{t}\cup K_{t}$ itself.
@@
We now prove Theorem 1.1, i.e., the unique determination of $\gamma$ from the DtN map $\Lambda_{\gamma}$. The proof is based on mathematical induction slice by slice.
4.1 Extract the solution space from DtN data
The only available data for the recovery oof the conductivity, aside from the graph structure, is the DtN matrix $\Lambda_{\gamma}$. In order to recover the conductivity using the boundary measurements from solutions in $\mathcal{U}^{(t)}$, we shall first prove that the respective boundary potentials $\ker T_{1}^{(t)}$ can be directly determined by $\Lambda_{\gamma}$. In fact, $T^{(t)}=\Lambda_{\gamma}(\partial D\setminus J_{t}^{\mathcal{S}};J_{t}^{\mathcal{S}})$ is a submatrix of $\Lambda_{\gamma}$, which achieves the unique determination by combining the following lemma:

### v2

@@ 2 Corner excitations
The set of localized potentials $\mathcal{U}^{(t)}$ is a linear subspace of $\mathbb{R}^{J_{t}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}}}$. The next result gives its complement. Essentially it is a rewriting of the governing equation in (1.1) for the localized solution in $\mathcal{U}^{(t)}$.
@@
Fix any $\mathbf{u}\in\mathcal{U}^{(t)}$ and extend it onto $D\cup\partial D$ by zero (still denoted by $\mathbf{u}$). By the definition of potential solutions, for $p\in L_{t+1}^{\mathcal{S}}$, we have ∑_q ∈N(p) γ_pq (u_q - u_p) = 0, or equivalently by $\operatorname{supp}(\mathbf{u})=L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}$, we have $\mathbf{v}_{p}\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}\cdot\mathbf{u}=0$. Thus we have
@@
Conversely, suppose that $\mathbf{u}\in(\operatorname{span}_{p\in L_{t+1}^{\mathcal{S}}}\{\mathbf{v}_{p}\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}\})^{\perp}$. Then,
$\mathbf{v}_{p}\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}\cdot\mathbf{u}=0,\quad\forall p\in L_{t+1}^{\mathcal{S}}.$
After extending $\mathbf{u}$ onto $D\cup\partial D$ by zero, the above equation implies that $\Delta_{\gamma}\mathbf{u}_{p}=0$ for all $p\in D$. Therefore, we have
$\mathbf{u}=S_{\gamma}(\mathbf{u}|_{\partial D})\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}\in\mathcal{U}^{(t)},$
and there holds $\mathcal{U}^{(t)}\supseteq(\operatorname{span}\{\mathbf{v}_{p}\big|_{L_{t}^{\mathcal{S}}\cup J_{t}^{\mathcal{S}}}:p\in L_{t+1}^{\mathcal{S}}\})^{\perp}$. Combining these two inclusions yields the desired orthogonal decomposition. ∎
@@
The definition of excitations and their properties do not rely on the explicit grid graph structure. Whenever one can divide the graph with the boundary nodes into two connected parts, one can define such localized spaces, and the respective properties remain valid.
@@ 3 Properties of lattice graphs
Note that the unique determination of the conductivity $\gamma$ from the DtN matrix $\Lambda_{\gamma}$ is not necessarily valid for a general graph structure. In fact, under the given definitions of $D$ and $\partial D$, if we choose the edge set to be complete in $D$ – that is, for every different nodes $p$ and $q$ in $D$, there exists an edge $pq$ in $E$ – then the number of undetermined conductivity parameters, scaling as $O(n^{2d})$, exceeds the size of DtN matrix $\Lambda_{\gamma}$, which grows as $O((2dn^{d-1})^{2})$, for large $n$. This naturally precludes the unique determination for graph structures, due to excessive connectivity (i.e. too many edge connections).
In this section, we identify several fundamental properties of lattice graphs that ensure the unique determination of the conductivity $\gamma$. For each integer $t$, consider the subgraph
@@
where the edge set $E^{\prime(t)}$ consists of all edges connecting nodes in $L_{t-1}^{\mathcal{S}}$ to nodes in $D\cup\partial D$. Associated to the subgraph $G^{\prime(t)}$, there exists a DtN map $\Lambda_{\gamma}^{\prime}$, acting on the boundary potentials supported on $J_{t-1}^{\mathcal{S}}\cup L_{t}$. Then we define the operator
@@
which maps the potentials prescribed on $L_{t}$ to the induced currents on $J_{t-1}^{\mathcal{S}}$.
The next lemma provides the key tool, which allows controlling the potential in the interior domain $L_{t}^{\mathcal{S}}$, using the potential and current both taken on $J_{t-1}^{\mathcal{S}}$. It can be viewed as the Cauchy problem for the discrete Laplacian $\Delta_{\gamma}$. Analogous to the Cauchy problem for the Laplacian, problem (3.2) is expected to be highly ill-conditioned.
@@
Let $d\leq t\leq dn$ and suppose that $\gamma>0$ is a conductivity on the lattice subgraph $G^{\prime(t)}$. The mixed boundary condition problem for $\mathbf{u}\in\mathbb{R}^{L_{t}^{\mathcal{S}}}$,
@@
has only the trivial solution, i.e., $\mathbf{u}=\mathbf{0}$ in $L_{t}^{\mathcal{S}}$.
@@
$H_{l}=(J_{t-1}^{\mathcal{S}}\cup L_{t}^{\mathcal{S}})\cap\{(x_{i})_{i}:x_{1}=l\},\quad H_{l}^{\mathcal{S}}=\bigcup_{i=0}^{l}H_{i}.$
Let $\mathbf{e}_{i}=(0,\ldots,0,{1},0,\ldots,0)\in\mathbb{R}^{d}$ be the $i$-th Cartesian coordinate vector. Note that for each $p\in H_{l}\setminus\partial D$ with $l\geq 1$, the node $q=p-\mathbf{e}_{1}$ is connected to exactly one node that lies outside $H_{l-1}^{\mathcal{S}}$, namely $p$ itself. The base case $\mathbf{u}|_{H_{0}}=\mathbf{0}$ holds by assumption. Now assume that $\mathbf{u}|_{H_{l-1}^{\mathcal{S}}}=\mathbf{0}$. For each $p\in H_{l}\setminus\partial D$, let $q=p-\mathbf{e}_{1}$. Then there holds
@@
The validity of Lemma 3.1 relies crucially on the grid graph structure. However, the next lemma is independent of the graph structure.
@@
For every integer $t$ with $d\leq t\leq dn$ and any strictly positive conductivity $\gamma>0$, the following three statements are equivalent:
@@
By interpreting the vector $\mathbf{c}$ as a potential supported on $G^{\prime(t)}$ with the zero boundary condition $\mathbf{c}|_{J_{t-1}^{\mathcal{S}}}=\mathbf{0}$, the relation (3.3) is equivalent to the uniqueness in the statement (i):
@@
The injectivity of $T_{2}^{\prime(t)}$ follows directly from Lemmas 3.2 and 3.1. The injectivity of the operator $T_{2}^{(t)}$ follows identically. ∎
@@
where the set $M_{p}:=\mathcal{N}(p)\cap\bigl(J_{t-1}^{\mathcal{S}}\cup L_{t-1}^{\mathcal{S}}\bigr)$ consists of the neighbors of $p$ within the preceding layers, and the edge set $E_{p}^{\prime\prime(t)}$ is given by
@@
Note that if the node $p$ is adjacent to the boundary, isolated boundary nodes may arise. The boundary currents on such isolated nodes are set to zero. Fig. 3 gives a schematic illustration of the removal of the node $p$ on the slice $t$. Note that the “hexagonal” appearance in Fig. 3 is due to the visualization from the diagonal angle of the underlying cubic lattice. This remark applies also to Figs. 4 and 5 below.
@@
Figure 3: Node deletion in dimension $d=3$. $L_{t}$ ((blue)), $L_{t+1}$ ((black)) and $K_{t}^{-}$ ((orange)) for $t\leq n+d-1$.
The following lemma establishes the unique recovery of the potential $\mathbf{u}$ on the reduced vertex set $L_{t}^{\mathcal{S}}\setminus\{p\}$ from the boundary measurements on $J_{t-1}^{\mathcal{S}}\setminus M_{p}$, i.e., the uniqueness of the interior potential recovery on the graph $G^{\prime\prime(t)}$.
@@
Let $d\leq t\leq dn$ and fix any node $p\in L_{t}$. Suppose that the conductivity $\gamma>0$ is specified on the subgraph $G_{p}^{\prime\prime(t)}$ defined in (3.4). Consider the boundary value problem:
$\left\{\begin{aligned} \Delta_{\gamma}\mathbf{u}&=\mathbf{0}\quad\text{in }L_{t-1}^{\mathcal{S}}\setminus M_{p},\\ \mathbf{u}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}},\\ D_{\gamma}\mathbf{u}&=\mathbf{0}\quad\text{on }J_{t-1}^{\mathcal{S}}\setminus M_{p}.\end{aligned}\right$. (3.5)
@@
Let the deleted node be located at $p=(p_{i})_{i=1}^{d}$. We can recover the potential $\mathbf{u}$ in the region $\{(x_{i})_{i=1}^{d}\in D:x_{1}<p_{1}\}$ using the argument of Lemma 3.1. By changing the inductive direction from $\mathbf{e}_{1}$ to $\mathbf{e}_{j}$ with $j>1$, $\mathbf{u}$ can also be recovered in the region
$\{(x_{i})_{i=1}^{d}:\exists i\in\{1,\ldots,d\}\text{ such that }x_{i}<p_{i}\}.$
Since these elements form the set $L_{t}^{\mathcal{S}}\setminus\{p\}$, we complete the recovery. ∎
This result yields an important relation among the vectors $\mathbf{v}_{q}$ associated with the nodes in the layers $L_{t}^{\mathcal{S}}$, given in the following corollary.
@@
The vectors $\mathbf{v}_{q}$ defined in (2.9) satisfy the following properties.
@@
    The collection of vectors $\{\mathbf{v}_{q}|_{(L_{t-1}^{\mathcal{S}}\cup J_{t-1}^{\mathcal{S}})\setminus M_{p}}\}_{q\in L_{t}^{\mathcal{S}}\setminus\{p\}}$ is linearly independent.
@@
    If $\mathbf{w}\in\operatorname{span}\{\mathbf{v}_{q}|_{L_{t-1}^{\mathcal{S}}\cup J_{t-1}^{\mathcal{S}}}\}_{q\in L_{t}^{\mathcal{S}}}$ is supported on $M_{p}$, then there exists $\alpha\in\mathbb{R}$ such that
    $\mathbf{w}=\alpha\mathbf{v}_{p}|_{L_{t-1}^{\mathcal{S}}\cup J_{t-1}^{\mathcal{S}}}.$
@@
The assertion (i) follows directly from the argument for the equivalence (i) $\Leftrightarrow$ (ii) in Lemma 3.2 and Lemma $3.$ To prove assertion (ii), suppose that the vector $\mathbf{w}=\sum_{q\in L_{t}^{\mathcal{S}}}c_{q}\mathbf{v}_{q}|_{L_{t-1}^{\mathcal{S}}\cup J_{t-1}^{\mathcal{S}}}$ is supported on $M_{p}$. By restricting $\mathbf{w}$ to $(L_{t-1}^{\mathcal{S}}\cup J_{t-1}^{\mathcal{S}})\setminus M_{p}$, we have
$\mathbf{0}=\sum_{q\in L_{t}^{\mathcal{S}}\setminus\{p\}}c_{q}\mathbf{v}_{q}|_{(L_{t-1}^{\mathcal{S}}\cup J_{t-1}^{\mathcal{S}})\setminus M_{p}}.$
By part (i), it follows that all coefficients $c_{q}=0$ for $q\neq p$. Hence, $\mathbf{w}=\alpha\mathbf{v}_{p}|_{L_{t-1}^{\mathcal{S}}\cup J_{t-1}^{\mathcal{S}}}$ for some $\alpha\in\mathbb{R}$. ∎
Finally, we state an elementary observation on interface connectivity.
@@
In particular, there are no edges connecting nodes within $L_{t}\cup K_{t}$ itself.
Remark 3.1.
The uniqueness analysis of this work essentially relies on the properties of the lattice graph structure listed in Proposition 3.1 and Lemmas 3.3 and 3.4, and the analysis strategy can be extended to slightly more general graphs, e.g., ball shaped bodies cut out of hypercubic lattices. However, the analysis of the case of other graphs, e.g., hexagonal grids, likely requires substantially new techniques and is still to be investigated.
@@
We now prove Theorem 1.1, i.e., the unique determination of $\gamma$ from the DtN matrix $\Lambda_{\gamma}$. The proof is based on mathematical induction slice by slice. The proof proceeds in the following two steps. (i) suppose that the conductivity $\gamma$ can be recovered on the edge set $E(L_{t-1}^{\mathcal{S}},\overline{D})$. Then we can determine the solution space $\mathcal{U}^{(t)}$. (ii) The solution space $\mathcal{U}^{(t)}$ yields a system of linear equations for $\gamma|_{E(L_{t}^{\mathcal{S}},\overline{D})\setminus E(L_{t-1}^{\mathcal{S}},\overline{D})}$, which is proven to be uniquely solvable in Theorem 4.1. Thus the conductivity $\gamma$ can be uniquely recovered on the enlarged edge set $E(L_{t}^{\mathcal{S}},\overline{D})$.
4.1 Extract the solution space from the DtN data $\Lambda_{\gamma}$
The only available data for the recovery of the conductivity $\gamma$, aside from the graph structure, is the DtN matrix $\Lambda_{\gamma}$. In order to recover the conductivity $\gamma$ using the boundary measurements from vectors in the solution space $\mathcal{U}^{(t)}$, we first prove that the respective boundary potentials in $\ker T_{1}^{(t)}$ can be directly determined by $\Lambda_{\gamma}$. In fact, $\ker T_{1}^{(t)}=\ker T^{(t)}$ , where $T^{(t)}=\Lambda_{\gamma}(\partial D\setminus J_{t}^{\mathcal{S}};J_{t}^{\mathcal{S}})$ is a submatrix of $\Lambda_{\gamma}$: